OpenStudy (anonymous):

To get a B in math, Alexandria must average 80 on five tests. Scores on the first four tests were 80, 74, 83, and 74. What is the lowest score that she can get on the last test and still get a B?

OpenStudy (anonymous):

So I have 80+74+83+74= 311

OpenStudy (anonymous):

311+x=400 or would it be 311+x=500

OpenStudy (anonymous):

It's certainly 80*5=400. All that's left is finding x.

OpenStudy (anonymous):

ok now I understand where the 400 comes from!

OpenStudy (anonymous):

Then x= 89

OpenStudy (anonymous):

Right.